OpenStudy (anonymous):

In triangle ABC, triangle C is a right angle. Find the remaining sides and angles. Round your answer to the nearest tenth. a=3, c=19

OpenStudy (raden):

i assumed that the remaining sides is b, so use the pythagorean theorem : b^2 = c^2 - a^2 b^2 = 19^2 - 3^2 b^2 = (19+3)(19-3) b^2 = 22 * 16 b = sqrt(22 * 16) = sqrt(19) * sqrt(16) = ....

OpenStudy (anonymous):

a.ma=9.1degrees, mb=80.9 degrees, b=18.8 b.ma=80.9 degrees, mb=9.1 degrees, b=18.8 c. ma=14.5 degrees, mb=75.5 degrees, b=18.8 d. ma=75.5 degrees, mb=14.5 degrees, b=18.8

OpenStudy (raden):

sqrt(19) = ... (use ur calculator) sqrt(16) = 4 multiply of them

OpenStudy (raden):

woops... sorry, i mean sqrt(22) * sqrt(16) = ...

OpenStudy (anonymous):

i got 18.761 is that right?

OpenStudy (raden):

yes, take nearst be 18.8

OpenStudy (raden):

now, to find the measure of angles A and be, u need trigonometry ...

OpenStudy (raden):

* A and B i mean :)

OpenStudy (anonymous):

How do i do that? sorry i am so lost